### 11 Dec 2023 v1

# Label Smoothing for Enhanced Text Sentiment Classification

Abstract
Label smoothing is a widely used technique in various domains, such as image classification and speech recognition, known for effectively combating model overfitting. However, there is few research on its application to text sentiment classification. To fill in the gap, this study investigates the implementation of label smoothing for sentiment classification by utilizing different levels of smoothing. The primary objective is to enhance sentiment classification accuracy by transforming discrete labels into smoothed label distributions. Through extensive experiments, we demonstrate the superior performance of label smoothing in text sentiment classification tasks across eight diverse datasets and deep learning architectures: TextCNN, BERT, and RoBERTa, under two learning schemes: training from scratch and fine-tuning.
**footnotetext: The corresponding author is Shijing Si, shijing.si@outlook.com.
1 Introduction
Text sentiment classification is to identify and extract emotional tendencies (such as positive, negative or neutral) from text, which provides important application scenarios and challenges for the development of natural language processing (NLP)(Li et al.,, 2022). Many methods have been proposed for this task, for instance, convolutional neural networks (CNN) (Zhang et al.,, 2015), and recurrent neural networks (RNN) with attention mechanism (Zhang et al.,, 2016; Chen et al.,, 2017). Transformer-based models (Jiang et al.,, 2021; Chang et al.,, 2020) have excelled in sentiment classification tasks, achieving state-of-the-art performance (Liu et al., 2022a, ; Liu et al., 2022c, ).
Label smoothing (LS) proceeds by using soft targets that are a weighted average of the hard targets and the uniform distribution over labels (Müller et al.,, 2019; Lienen and Hüllermeier,, 2021). This technique lessens the disparity between the top probability estimate and the remaining ones, thereby acting as a barrier to the model from generating extremely confident predictions, consequently decreasing the model’s likelihood of becoming excessively tailored to the training data (Lukasik et al.,, 2020; Gao et al.,, 2020). This also fosters generalization, leading to better performance on unseen data (Chen et al.,, 2020; Cui et al.,, 2022). Furthermore, LS can also mitigate the impact of noisy labels on the training process(Li et al.,, 2020).
LS has achieved widespread success in NLP (Chandrasegaran et al.,, 2022; Liu et al., 2022b, ) by introducing soft targets, which allows the model to optimize towards a more flexible direction during training(Haque et al.,, 2023; Pan et al.,, 2022). However, the research on LS for text sentiment classification is still limited. For the task of sentiment classification, emotion is not an absolute and discrete concept (Pang et al.,, 2002; Onan,, 2022), but there is certain fuzziness and continuity (Hasib et al.,, 2021). LS can take account of this ambiguity, allowing the model to learn more balanced relative probabilities between different emotions.
In order to validate whether LS can benefit sentiment classification, we conduct extensive experiments on three widely used deep neural network architectures: TextCNN, BERT, and RoBERTa, under two learning schemes: training from scratch and fine-tuning. We investigate the performance of LS on six three-class datasets and two binary-class datasets. A comparison with the baseline models is also conducted. By varying the smoothing parameter, LS approach can outperform the baseline method across all architectures.
The remaining sections of this paper are organized as follows. We first review the related work in Section 2. Then, in Section 3, we propose the application of LS in text sentiment classification, including the definition of the text LS learning problem, the theoretical foundations of text LS learning and the selected deep learning architecture. Subsequently, in Section 4, we present the experimental results. Finally, Section 5 provides a discussion, and Section 6 concludes the paper.
2 Related Work
Text sentiment classification is one special type of text classification Si et al., (2020); Kowsari et al., (2019), as the labels are ordinal Si et al., (2022). Since the 2010s, deep learning models (Zulqarnain et al.,, 2020; Liu et al.,, 2016), which automatically extract meaningful representations for text without the need for manual rule and feature design(Minaee et al.,, 2021), are increasingly utilized for text classification.
LS is firstly introduced in the field of computer vision (CV) and has achieved success in various visual recognition tasks(Gao et al.,, 2017; Müller et al.,, 2019). Later, the method is shown to be effective in Machine Translation(MT)(Gao et al.,, 2020). Moreover, it also has applications in sentiment classification(Luo et al.,, 2022, 2021) and Named Entity Recognition(NER)(Zhu and Li,, 2022; Yu et al.,, 2022), improving model calibration and bringing flatter neural minima.
Meanwhile, researchers have started to apply LS to text sentiment classification. One attempt is to add LS to the loss function to enhance the performance of the model (Wang et al., 2022a, ). Similarly, another group add LS to loss function and perform emotion classification on the adaptive fusion features obtained (Yan et al.,, 2023). Also, properties of LS and its adversarial variants are studied, showing LS can enhance the adversarial robustness of the model(Yang et al.,, 2022). Yan et al., (2023) proposes a new cyclic smoothing labeling technique for handling the periodicity of angles and increasing error tolerance for adjacent angles. They also designed a densely-coded tag that greatly reduced the length of the code. Wu et al., (2022) propose an efficient data augmentation method, termed text smoothing, by converting a sentence from its one-hot representation to a controllable smoothed representation and show that text smoothing outperforms various mainstream data augmentation methods by a substantial margin.
Although LS is a common label processing technique, there are still some problems and challenges that need further study. Examples include comparison of effects across different data sets, tasks, and model structures, comparison with other regularization methods, etc(Huang et al.,, 2020). Studying LS can explore these problems and promote the further development of the field of LS. In this paper, we will focus on how to exploit LS in text classification, and a discussion about its superiority and principle will also be given.
3 LS Method on Texts Classification
3.1 Label Smooth Processing
Given a tokenized input $(x_{1},x_{2},x_{3},\ldots,x_{n})$ and a set of labels $(y_{1},y_{2},y_{3},\ldots,y_{k})$, where $n$ is the length of input and $k$ is the number of categories for classification. Thus, $D_{i}=(d_{x_{i}}^{y_{1}},d_{x_{i}}^{y_{2}},d_{x_{i}}^{y_{3}},\ldots,d_{x_{i}}^{y_{k}})$ describes the extent to which the instance $x_{i}$ belongs to a label $y_{k}$. We expect to obtain a final classification result $d_{i}$ which is typically the maximum value of the label distribution $D_{i}$ obtained by applying a softmax normalization to the last layer of the deep neural network. The label corresponding to the maximum value represents the category to which this text segment belongs.
Figure 1: Example of single-label, multi-label and LS
Take the instance in Figure 1 as an example, in single-label conditions, $d_{x}^{y_{2}}=1$ and the rest labels all equal to 00. In multi-label conditions, $d_{x}^{y_{2}}=d_{x}^{y_{5}}=0.5$ and the rest labels are 00. Both two methods only represents part of the labels and ignore the others. Finally, in LS, all of the labels are given concentrations and they also add up to $1$. Meanwhile, the dominance of The most significant label $d_{x}^{y_{2}}$ should be maintained to ensure that the model learns the primary features for sentiment classification.
Due to the continuity of label distribution, it has a larger labeling space, a broader range of expression, and therefore can provide greater flexibility in the learning process. Specifically, in order to convert hard labels to soft label distributions, we use LS method, performing a linear transformation on the original label distribution $D_{i}=(d_{x_{i}}^{y_{1}},d_{x_{i}}^{y_{2}},d_{x_{i}}^{y_{3}},\ldots,d_{x_{i}}^{y_{k}})$ to obtain $D_{i}^{{}^{\prime}}=(1-k\lambda)D+\lambda,\lambda\in[0,1]$.
This method also addresses the need for manual labeling. One of the primary challenges in the task of LS is the difficulty of obtaining the true label distribution. Most classification datasets do not provide this information, and theoretically, acquiring the precise label distribution would require extensive manual labeling of the same sample to obtain its statistical distribution, which is prohibitively expensive.
Suppose that $\theta$ is the parameters of the models waiting to be trained, and $P(y_{k}|x_{i};\theta)=p(y_{1}|x_{i};\theta),p(y_{2}|x_{i};\theta),p(y_{3}|x_{i};\theta),\ldots,p(y_{k}|x_{i};\theta)$ is the probability distribution finally output by the model. We use Kullback-Leibler divergence in model’s training process, the reason and its difference with cross entropy will be discussed thoroughly in Section 4.2. Our goal is to find best parameters $\theta^{*}$ that satisfies:
$\begin{split}\theta^{*}&=\arg\min\limits_{\theta}\sum_{i}KL(D^{{}^{\prime}}_{i}||P_{i})\\ &=\arg\min_{\theta}\sum_{i}\sum_{j}\left(d^{y_{k}}_{x_{i}}\ln\frac{d^{y_{k}}_{x_{i}}}{p(y_{k}|x_{i};\theta)}\right)\end{split}$ (1)
This expression can be transformed into a maximum likelihood function:
$\theta^{*}=\arg\max_{\theta}\sum_{i}\sum_{j}\left(d^{y_{k}}_{x_{i}}\ln p(y_{k}|x_{i};\theta)\right)$ (2)
In the meantime, we’ll look at the optimization equation in traditional single-label and multi-label methods, they use cross entropy as the main practice:
$\theta^{*}=H(D^{{}^{\prime}}_{i},P_{i})=-\sum_{i}\sum_{k}d^{y_{k}}_{x_{i}}\log p(y_{k}|x_{i};\theta)$ (3)
$\begin{split}\theta^{*}&=\sum_{i}\frac{1}{\lvert Y_{i}\rvert}H(D^{{}^{\prime}}_{i},P_{i})\\ &=-\sum_{i}\sum_{k}\frac{1}{\lvert Y_{i}\rvert}d^{y_{k}}_{x_{i}}\log p(y_{k}|x_{i};\theta)\end{split}$ (4)
Eq.3 and Eq.4 are the optimization functions of single-label and multi-label method respectively. Their corresponding maximum likelihood functions are as Eq.5 and Eq.6:
$\theta^{*}=\arg\max_{\theta}\sum_{i}\ln p(y(x_{i})|x_{i};\theta)$ (5)
$\theta^{*}=\arg\max_{\theta}\sum_{i}\frac{1}{|Y_{i}|}\sum_{y\in Y_{i}}\ln p(y|x_{i};\theta)$ (6)
For single-label problems, where the model only considers one label, it is necessary to compute the entropy of that particular label. In contrast, multi-label tasks require aggregating the entropies of multiple labels. Compare Eq.2, Eq.5 and Eq.6, the simplified maximum likelihood functions differ only from the ranges of labels involved in the calculation(Geng,, 2016). ). Single- label and multi-label learning are special cases of the distribution of LS. LS can be seen as a more general learning technique due to its correlationship with traditional training methods.
3.2 Training Deep learning Models
Now we need to train the deep learning model. The goal of model training is to iteratively optimize the parameter $\theta$ in order to find the optimal parameter $\theta^{*}$ that satisfies:
$\arg\min\limits_{\theta}\sum_{i}KL(D^{{}^{\prime}}_{i}||P_{i})$ (7)
Thus the loss function can be defined as:
$T=-\sum_{i}KL(D^{{}^{\prime}}_{i}||P_{i})$ (8)
In the forward propagation, there is:
$output=forward(input,\theta_{i})$ (9)
Where $input$ is the input data $(x_{i1},x_{i2},x_{i3},\ldots,x_{ik})$, $\theta$ is the parameter value at time step $i$, $output$ is the result of the model. The specific calculation process of forward propagation varies according to the structure of the model.
Then, We use SGD method to optimize the model. In the backward propagation, according to the chain rule , we have:
$\delta_{\text{output}}=\frac{\partial\text{T}}{\partial output}$ (10)
$\delta_{i}=\delta_{i+1}\cdot\frac{\partial a_{i}}{\partial z_{i+1}}\cdot\frac{\partial z_{i}}{\partial a_{i}}$ (11)
For each layer $i$, the gradient is propagated backward from the output layer. $\delta_{i}$ represents the gradient of layer $i$, $\delta_{i+1}$ represents the gradient of the next layer, the input for each layer is $z_{i}$, the output of the activation function for each layer is $a_{i}$.
The SGD algorithm randomly selects a group from the sample, updates it and then extracts another group and updates it again(Adikari and Draper,, 2021). In the case of extremely large sample size, it is possible to obtain a model with a loss value within the acceptable range without training all samples.(Prasher et al.,, 2023; Subramaniam et al.,, 2023) The iteration update formula for SGD (Stochastic Gradient Descent) is:
$\theta_{i+1}=\theta_{i}-\eta\cdot\delta_{i}$ (12)
where $\theta_{i+1}$ is the parameter vector at iteration $i+1$, $\theta_{i}$ is the parameter vector at iteration $i$, $\eta$ is the learning rate, which controls the step size of each update. $\delta_{i}$ is the gradient of the loss function T with respect to the parameter $\theta_{i}$.
3.3 Selected Deep Dearning Architectures
TextCNN, BERT, and RoBERTa are three highly influential and successful deep learning architectures in the field.
TextCNN, as a pioneering architecture, introduced the idea of applying convolutional neural networks to text classification tasks. It leverages convolutional layers with various kernel sizes to capture local features in text data. TextCNN has demonstrated strong performance and efficiency in handling fixed-length input sequences(Wang et al., 2022b, ; Jiang,, 2022).
Further, BERT (Bidirectional Encoder Representations from Transformers) brought about a breakthrough in language understanding. BERT introduced the concept of pre-training a transformer-based model on large amounts of unlabeled text data and fine-tuning it for downstream tasks. It showed remarkable success in a wide range of natural language processing tasks, including question answering, sentiment analysis, and named entity recognition.
RoBERTa, building upon BERT’s foundation, refined the pre-training process and achieved improved performance. It incorporated additional pre-training data and introduced modifications to the training objectives. RoBERTa demonstrated enhanced language representation capabilities and surpassed BERT’s performance on various benchmarks and tasks.
In our experiments, we aim to test the generalizability of LS across different architectures. By evaluating our method on these three models, we aim to showcase the applicability and effectiveness of this approach across diverse deep learning architectures.
(I) TEXTCNN
The structure of TextCNN consists of a convolutional layer, a max pooling layer, and a softmax layer for n-class classification.
Figure 2: Workflow of TextCNN sentiment classification architecture
Firstly, the textual input will undergo convolutional operations using different sizes of convolutional kernels, aiming to capture the textual features(Zhang et al.,, 2020). Assuming the input is a text sequence $x=(x_{i1},x_{i2},x_{i3},\ldots,x_{ik})$, where $x_{i}$ represents the word vector for the $i-th$ word. The convolutional layer applies multiple convolutional filters to extract local features:
$c_{i}=f(W\cdot x_{i:i+h-1}+b)$ (13)
Here, $x_{i:i+h-1}$ represents the concatenation of word vectors from the $i-th$ word to the $(i+h-1)-th$ word, W is the weight matrix of the convolutional filters, b is the bias vector, and $f$ denotes the activation function.
After the convolution is completed, the output will undergo a pooling layer for further processing. The max pooling operation selects the maximum value from each feature map as the pooled feature value:
$p_{i}=\max(c_{i:i+h-1})$ (14)
Here, $c_{i:i+h-1}$ represents the feature vector from the $i-th$ feature to the $(i+h-1)-th$ feature.
Finally, The pooled feature vector serves as the input to the fully connected layer, which performs linear transformation and non-linear activation,:
$y=\text{softmax}(W_{y}\cdot p+b_{y})$ (15)
$W_{y}$ is the weight matrix of the fully connected layer, $b_{y}$ is the bias vector, and $softmax$ is the activation function for multi-class classification.
TextCNN’s simplicity, lower parameter count, and computational efficiency contribute to its fast training speed(Suyan et al.,, 2019). It has demonstrated good performance in text classification tasks.
(II) BERT
BERT (Bidirectional Encoder Representations from Transformers) is a pre-trained language model that utilizes a transformer-based architecture for various natural language processing tasks(Palani et al.,, 2021). The model incorporates both left and right context during training, allowing it to capture deep contextual relationships in text(Sebastian et al.,, 2022).
Figure 3: Workflow of BERT sentiment classification architecture
The input to BERT consists of tokenized text, where each token represents a word. Special tokens like [CLS] and [SEP] are added at the beginning and end of each sentence respectively. Let $H$ be the input sequence $X$ processed by the embedding layer.
Input $H$ is to be split into multiple heads: $H=[h_{1},h_{2},\ldots,h_{h}]$, where each $h$ is a head containing a subset of hidden states.
For each head $h$, we initialize its own set of weight matrix: $Key(K)$, $Value(V)$, and $Query(Q)$. These heads are then linearly transformed and passed through the softmax function to obtain:
$\begin{split}Z_{i}&=\text{Attention}(QW^{Q}_{i},KW^{K}_{i},VW^{V}_{i})\\ &=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V\end{split}$ (16)
The outputs of each head are concatenated to form the multi-head attention output:
$\begin{split}Z&=\text{MultiHead}(Q,K,V)\\ &=\text{Concat}(Z_{1},Z_{2},\ldots,Z_{h})\end{split}$ (17)
This step involves repeating the self-attention operation multiple times, allowing the model to learn information from different perspectives and enrich the semantic representation.
Next, the result $Z=[z_{1},z_{2},\ldots,z_{h}]$ is fed into a feed-forward neural network layer. This layer applies multiple matrix transformations and activation functions introducing non-linearity and obtaining the final vector representation:
$F=relu(W_{2}\cdot(W_{1}Z+b_{1})+b_{2})$ (18)
After the feed-forward neural network layer, a pooling operation is performed to obtain the aggregated representation. The pooling method used in this experiment is max pooling, where the maximum value is selected across each dimension of the representation.
Then, the pooled representation is passed to the classification layer. This layer typically consists of one or more fully connected layers followed by a softmax activation function. The classification layer applies a series of linear transformations and activation functions to the pooled representation, mapping it to corresponding class probabilities.
BERT employs a stack of transformer encoders, composed of self-attention and feed-forward neural network layers. This allows the model to attend to different parts of the input sequence and capture word dependencies. BERT is trained using Masked Language Model(MLM), where a percentage of input tokens are masked, and the model predicts these masked tokens based on context. It also uses Next Sentence Prediction (NSP), where it learns to predict if a second sentence follows the first(Trang and Shcherbakov,, 2020). BERT can be fine-tuned on specific tasks by adding task-specific layers and training on labeled data(Riaz et al.,, 2022). BERT has achieved state-of-the-art results in various NLP tasks, demonstrating its effectiveness in understanding context and semantic relationships in text(Zou et al.,, 2022).
(III) RoBERTa
The RoBERTa model is an improved version built upon the BERT pre-trained model. RoBERTa introduces three key enhancements to the BERT model(Kamath et al.,, 2022). Firstly, it adopts a dynamic masking mechanism, which generates a new masking pattern for each input sequence presented to the model. This allows for greater variation and exploration during training. Secondly, RoBERTa removes the Next Sentence Prediction task, which was present in BERT. By eliminating NSP, the model’s efficiency is improved as it no longer needs to predict the relationship between two sentences(Xu and Hou,, 2022; Choure et al.,, 2022). Lastly, RoBERTa increases the size of the training corpus and adjusts the batch size and training steps accordingly(Mu et al.,, 2021; Qi et al.,, 2022). This expansion of training data and adjustments in batch size and steps contribute to enhanced learning and better model performance.
4 Experimental Results
We construct five LS prediction models incorporating varying degrees of LS and assess their performance compared to a baseline model on eight data sets using both BERT, TEXTCNN and RoBERTa architectures. The experiments are conducted to evaluate the effectiveness of LS methods in improving the prediction accuracy and robustness of the models.
4.1 Data Sets
We utilize publicly available data sets obtained from Kaggle and Hugging Face for our study. The data set collection comprises eight distinct data sets, of which six are three-class and two are two-class. The diverse range of data sets enables us to evaluate the performance and generalizability of our models across various data types and classification scenarios.
Twitter Financial News Sentiment(TFNS) was collected using the Twitter API. It holds 11,932 documents annotated with 3 labels and comprises annotated finance-related tweets.
Kaggle Financial Sentiment(KFS) is a dataset of financial reviews annotated with 3 labels. These reviews cover topics such as stocks, investments, market analysis, corporate performance, technology, and real estate.
Tweet Sentiment Extraction(TSE) is a three-class dataset from a Kaggle contest. It has 27481 samples of a tweet and a sentiment label.
Auditor Sentiment(AS) gathers the auditor evaluations into one dataset. It contains thousands of sentences from English financial news, grouped into three categories by emotion.
Financial Phrasebank(FP) consists of 4840 sentences from English language financial news categorised by sentiment. The dataset is divided into three classes by agreement rate of 5-8 annotators.
Chatgpt Sentiment Analysis(CSA) contains 10,000 pieces of data. The dataset gathers tweets about Chatgpt and Reflects people’s perception of it. Emotions are divided into three categories.
Rotten Tomatoes Reviews(RTR) consists of 10,662 processed sentences from Rotten Tomatoes movie reviews. It’s a two-class dataset with 5,331 positive and 5,331 negative.
Sentiment140(Sent140) is a widely used sentiment analysis dataset created by researchers at Stanford University. It contains 1,600,000 tweets extracted using the twitter api. The tweets have been annotated to two classes. In this work, we only use parts of the data.
4.2 Models and Loss Functions
Models: Table 1 presents the configuration of the models and loss functions used in the experiment. The LS models (LS1-LS5) employ varying levels of LS, with the smoothing intensity increasing from LS1 to LS5. The original labels are used as they are for LS1, while for LS2 to LS5, the labels are smoothed by adding small perturbations. By gradually increasing the smoothing level from LS1 to LS5, we explore the effects of different degrees of LS on the model’s accuracy and stability. All LS models utilize the KL Loss as the loss function. This loss function measures the divergence between the predicted probability distribution and the smoothed label distribution. In contrast, the baseline model uses single labels without any smoothing and employs the Cross Entropy loss function.
Table 1: Model smoothing levels and loss functions. The original one-hot label is $\mathbf{y}$, and the soft label $\mathbf{y}^{LS}$
Model Smoothing Level (three-class) Smoothing Level (two-class) Loss Function
LS2 $\mathbf{y}^{LS}=\mathbf{y}*0.97+0.01*\mathbf{1}$ $\mathbf{y}^{LS}=\mathbf{y}*0.98+0.01*\mathbf{1}$ KL Loss
LS3 $\mathbf{y}^{LS}=\mathbf{y}*0.925+0.025*\mathbf{1}$ $\mathbf{y}^{LS}=\mathbf{y}*0.9+0.05*\mathbf{1}$ KL Loss
LS4 $\mathbf{y}^{LS}=\mathbf{y}*0.85+0.05*\mathbf{1}$ $\mathbf{y}^{LS}=\mathbf{y}*0.8+0.1*\mathbf{1}$ KL Loss
LS5 $\mathbf{y}^{LS}=\mathbf{y}*0.7+0.1*\mathbf{1}$ $\mathbf{y}^{LS}=\mathbf{y}*0.7+0.15*\mathbf{1}$ KL Loss
Baseline $\mathbf{y}$ $\mathbf{y}$ Cross Entropy
Cross-Entropy: In a regular multi-classification task, cross-entropy is used to measure the difference between the model’s predicted results and the true labels. Assume there are classes, and each sample’s true label can be represented as a n-dimensional one-hot encoded vector p, where only one element is 1, indicating the corresponding class to which the sample belongs. And assume the model’s predicted result is also a n-dimensional probability distribution vector q, where each element represents the probability of the sample belonging to the corresponding class. The definition of the cross-entropy loss function H(p,q) is as follows:
$H(p,q)=-\sum_{i=1}^{n}p(x_{i})\log q(x_{i})$ (19)
The cross-entropy loss function calculates the negative logarithm of the product of the true label probability and the predicted probability for each class, summed over all classes. It penalizes the model for predicting low probabilities(Udhayakumar et al.,, 2019). Minimizing the cross-entropy loss function helps the model make more accurate predictions of the sample’s class.
KL Divergence: KL divergence measures the information difference between two probability distributions. In LS, it assesses the dissimilarity from the ground truth distribution to the predicted distribution. It emphasizes the discrepancy in information when transitioning from the true distribution to the predicted distribution. The formula is shown as follows:
$KL(p||q)=\sum_{i=1}^{n}p(x_{i})\log\left(\frac{p(x_{i})}{q(x_{i})}\right)$ (20)
Thus, we have:
$\begin{split}KL(p||q)=&-\sum_{i=1}^{n}p(x_{i})\log q(x_{i})\\ &+\sum_{i=1}^{n}p(x_{i})\log p(x_{i})\end{split}$ (21)
$KL(p||q)=H(p,q)-H(p)$ (22)
Although the formulas of cross-entropy and KL divergence differ only by the inclusion of the entropy of the true labels, their distinct implications and applications justify their separate usage(Chai et al.,, 2017). In LS, KL divergence is commonly preferred as an objective function due to the following reasons:
(I)Sensitivity to Distribution Shift: KL divergence is highly sensitive to shifts between distributions, making it suitable for capturing the accuracy of predicted distributions in reflecting the characteristics of the ground truth distribution, beyond just class labels.
(II)Handling Asymmetry: In certain scenarios, the true and predicted distributions may exhibit asymmetry, with varying degrees of dissimilarity. KL divergence handles this asymmetry better as it focuses on the information difference from the true distribution to the predicted distribution.
4.3 Results
The results produced by BERT, TEXTCNN and RoBERTa architectures are presented in Table 2. The accuracy curves are presented in Figure 4, Figure 5 and Figure 6 respectively, depicting the performance of the models at each epoch.
On the BERT architecture, LS models consistently outperform the baseline model across multiple datasets. Specifically, LS2 achieves the highest accuracy in most cases, closely followed by LS4 and LS5. This indicates that incorporating LS with varying degrees of LS can significantly improve classification accuracy, particularly for binary and ternary classification datasets.
Moving to the TextCNN architecture, LS models exhibit superior performance on ternary classification datasets compared to binary classification datasets. LS2 and LS4 consistently outperform the other models in terms of accuracy. This highlights the effectiveness of LS when implemented in the TextCNN architecture for enhancing classification outcomes. Although the training acceleration of LS models on the CNN architecture is slower, they have greater potential in the later epochs.
On the RoBERTa architecture, LS models demonstrate competitive performance, LS2 and LS5 achieving the highest accuracy in most cases. This suggests that LS can effectively leverage the strengths of the RoBERTa architecture to achieve improved classification accuracy.
The experimental results demonstrate that LS models, when integrated into different architectures, consistently outperform the baseline model. LS2, LS4, and LS5 consistently demonstrate superior accuracy, indicating the efficacy of LS in enhancing classification performance. These findings emphasize the potential of LS as a valuable approach for improving classification outcomes in various tasks and architectures.
The results emphasize the effectiveness of the LS approach in improving the accuracy of the models. The varying degrees of LS in the LS models provide a means to balance model sensitivity to noisy labels and mitigate overconfidence. This contributes to the models’ enhanced robustness and generalization capabilities, ultimately leading to improved performance in sentiment classification tasks.
Table 2: VALIDATION ACCURACY ON DIFFERENT ARCHITECTURES
Architecture Algorithm TFNS KFS TSE AS FP CSA RTR Sent140
BERT LS2 0.8769 0.7981 0.7917 0.8411 0.8930 0.7202 0.7820 0.7636
LS3 0.8744 0.7964 0.7912 0.8318 0.8955 0.7181 0.7780 0.7647
LS4 0.8710 0.7913 0.7881 0.8390 0.8968 0.7254 0.7780 0.7711
LS5 0.8740 0.7956 0.7889 0.8421 0.8955 0.7181 0.7860 0.7594
Baseline 0.8689 0.7896 0.7881 0.8256 0.8803 0.7149 0.7780 0.7626
TextCNN LS2 0.8237 0.6852 0.7074 0.7761 0.8382 0.7234 0.7860 0.7455
LS3 0.8233 0.6801 0.7029 0.7740 0.8357 0.7245 0.7920 0.7540
LS4 0.8216 0.6792 0.7066 0.7771 0.8369 0.7183 0.7780 0.7455
LS5 0.8204 0.6843 0.7029 0.7730 0.8344 0.7338 0.7860 0.7444
Baseline 0.8141 0.6852 0.7049 0.7647 0.8319 0.7090 0.7920 0.7626
RoBERTa LS2 0.8915 0.8375 0.7946 0.8689 0.9031 0.8101 0.8680 0.8588
LS3 0.8961 0.8263 0.8036 0.8617 0.9070 0.8029 0.8660 0.8524
LS4 0.8957 0.8246 0.7999 0.8658 0.9108 0.8163 0.8720 0.8610
LS5 0.8978 0.8298 0.8024 0.8638 0.9045 0.8122 0.8780 0.8706
Baseline 0.8769 0.8315 0.7991 0.8566 0.8994 0.8039 0.8720 0.8759
Figure 4: Accuracy on the validation set at different epochs using BERT
Figure 5: Accuracy on the validation set at different epochs using TEXTCNN
Figure 6: Accuracy on the validation set at different epochs using RoBERTa
Figure 7: Visualization of Dimension Reduction of the BERT Model
5 Discussions
In this section, we attempt to analyze why text LS can achieve higher accuracy under the same data and parameter conditions.
Higher Accuracy: LS methods can achieve higher accuracy. The impact of label distribution can be profound, as it effectively addresses certain challenges in sentiment classification. Label distribution helps alleviate overconfidence in model predictions and reduces sensitivity to noisy labels, resulting in improved robustness and generalization performance.
Furthermore, in sentiment classification, emotions are not always purely positive or negative. There can be instances where a dominant emotion encompasses other subtle emotions. LS models excel at capturing and representing such phenomena, allowing for a more nuanced understanding and interpretation of sentiment.By considering a broader spectrum of emotions and incorporating LS techniques, LS models offer a more comprehensive and accurate representation of sentiment classification, leading to improved performance in sentiment analysis tasks.
Acceleration of Convergence: We further analyzed the convergence performance of LS on the BERT, TextCNN, and RoBERTa architectures. It can be observed that in many cases, LS models implemented on BERT and RoBERTa exhibit significantly faster convergence compared to the baseline model, and can even achieve relatively high accuracy within the initial few epochs of training, as shown in Figure 4 and Figure 6. This also indicates that LS is capable of effectively saving computational resources, reducing the training time, and improving overall efficiency.
Robust Performance: The LS methods implemented on the three deep learning architectures consistently achieve stable and superior prediction performance on most of the datasets, demonstrating the robustness of the models. This indicates that the LS approach is effective in handling diverse data scenarios and can generalize well to unseen examples. Furthermore, the ability of the models to consistently outperform the baseline methods across multiple datasets highlights their reliability and suitability for real-world applications. The robust predictive performance of LS methods reinforces their value and reinforces their potential as a powerful tool in various domains requiring accurate and reliable predictions.
Reduce Overfitting: LS and KL divergence can help reduce overfitting compared to cross-entropy for several reasons. Cross-entropy loss assigns high confidence to a single target class and penalizes all other classes. This can lead to overconfident predictions, especially when the training data is limited or imbalanced. In contrast, LS considers the entire label distribution, allowing the model to capture the relationships between classes and reduce the risk of overfitting to individual examples.
LS: LS further mitigates overfitting. The effects of LS can be reflected in the hidden layer of the Model. Figure 7 visualizes similarities and differences between data points by mapping high-dimensional features into two - dimensional spaces and depicts the last layer of BERT models with different smoothing levels. The data points in single-label model overlap, while model with smoothing labels can distinguish data points well. After applying the LS method, during the training phase, the loss doesn’t decrease too rapidly when the predictions are correct, and it doesn’t penalize too heavily when the predictions are wrong. This prevents the network from easily getting stuck in local optima and helps to mitigate overfitting to some extent. Additionally, in scenarios where the classification categories are closely related, the network’s predictions are not excessively absolute.
6 Conclusion
In this paper, we delve into the application of LS in text sentiment classification, a key area of Natural Language Processing (NLP). Our experiments demonstrate that by fine-tuning the appropriate smoothing parameters, LS can achieve improved prediction accuracy across multiple deep learning architectures and various test datasets. However, given the inherent emotional bias in different texts, constructing more precise textual sentiment distribution labels to capture subtle variations in human emotions remains a challenging task. This highlights the need for further research and development in this area to fully harness the potential of LS in enhancing the performance and robustness of sentiment classification models.